# PartSAM: A Scalable Promptable Part Segmentation Model Trained on Native 3D Data

### v1

@@ A.2.2 Ablation of Training Data
A.2.3 Application
@@
Amodal Part Segmentation AI-generated meshes are often presented as a unified whole, with initial segmentation typically resulting in fragmented parts. Due to PartSAM’s strong generalization capabilities, it can accurately segment complex structures even when they are incomplete or fragmented. By integrating PartSAM with the part completion model HoloPart (Yang et al., 2025), these segmented fragments can be seamlessly reconstructed, achieving a comprehensive and complete decomposition of the mesh. This combined approach ensures that the segmented parts are also fully realized, overcoming the limitations of standard segmentation techniques.
A.2.4 Results on PartObjaverse-Tiny with Mesh Connectivity
In the evaluation of automatic segmentation on PartObjaverse-Tiny (Section 4.2), we compare PartField (Liu et al., 2025) and our method in a setting where mesh connectivity is disabled. The rationale behind this choice is further explained through the quantitative results presented in Table 4. Rather than training PartField, we use a randomly initialized model to generate the feature field and apply agglomerative clustering to produce segmentation results with mesh connectivity (PartField (B) in Table 4). Interestingly, the segmentation performance, in terms of IoU, is significantly better than that of the pre-trained PartField when mesh connectivity is disabled (PartField (A)), with a $27\%$ improvement.
Moreover, as indicated by the results in the last row of Table 4, we find that PartSAM can also benefit from mesh connectivity information. By applying graph cuts (Boykov et al., 2001) to the segmentation results for one iteration, PartSAM outperforms PartField under this setting (Ours (B)), further demonstrating the flexibility of our method.
A.2.5 Complexity Analysis
@@
We present the complexity analysis in Table 5, reporting both the inference time on a single NVIDIA H20 GPU and the number of network parameters. Compared with lifting-based methods that require per-shape optimization (Tang et al., 2024; Yang et al., 2024b), PartSAM achieves substantially faster inference due to its feed-forward design. Compared with PartField (Liu et al., 2025), PartSAM incurs only a slight increase in inference time for segmenting an entire shape. Although our encoder is built upon PartField, it eliminates the need for per-face dense feature sampling and clustering. Furthermore, our decoder is lightweight, ensuring that the overall inference time remains well-controlled. These results demonstrate that our method effectively balances computational cost and segmentation performance.
@@
Methods Time Params
SAMesh $\sim$ 7min /
SAMPart3D $\sim$ 15min 114M
PartField $\sim$ 10s 106M
Ours $\sim$ 12s 118M
A.2.6 Limitations.
Although PartSAM achieves state-of-the-art performance in class-agnostic part segmentation, similar to 2D SAM (Kirillov et al., 2023; Ravi et al., ), it cannot directly produce semantic labels for these masks, which are important for some downstream tasks. One possible future direction to address this limitation is to generate large-scale training datasets consisting of 3D shapes, part annotations, and corresponding semantic labels. This could potentially be achieved by utilizing PartSAM’s interactive segmentation capabilities, where users provide feedback on segmented parts, facilitating the assignment of semantic categories.
A.3 Additional Visualization
@@
Visualization of Segmentation Results Additional segmentation results of PartSAM are shown in Figure 11, where it consistently delivers impressive performance across different kinds of shapes.

### v3

@@ A.2.2 Ablation of Training Data
A.2.3 Advantage of 3D Native Data and Scaling Analysis
In Figure 10, we compare our native 3D part labels with the 2D-lifting results produced via community detection in Tang et al. (2024). The lifted 2D masks are frequently incomplete and overly fragmented due to heavy occlusions, making them unsuitable as reliable supervision in our framework. In contrast, our native 3D annotations maintain strong spatial coherence and accurately capture interior structures. Notably, even when the outer clothing of a character model is removed (second column), our 3D labels still reveal clean and consistent part segmentation beneath the surface—an outcome fundamentally unattainable for 2D-lifting methods that only observe the visible exterior. The advantage of precise interior part segmentation is also illustrated in Figure 9.
To further evaluate the effect of large-scale native 3D training, in Figure 11, we present the scaling curve of PartSAM, showing how segmentation performance improves with increasing training data size, from 40k to 500k shapes. The results indicate a consistent improvement in both interactive and automatic segmentation, suggesting that PartSAM continues to benefit from large-scale 3D-native supervision. Notably, the performance gains persist even at the largest data size, demonstrating that the model is scalable and benefits from large-scale 3D-native supervision.
Figure 11: Scaling curve of PartSAM with respect to training data size. The plot shows segmentation accuracy (IoU for interactive segmentation with 1 prompt point and automatic segmentation) as the training data increases from 40k to 500k shapes.
A.2.4 Application
Amodal Part Segmentation AI-generated meshes are often presented as a unified whole, with initial segmentation typically resulting in fragmented parts. Due to PartSAM’s strong generalization capabilities, it can accurately segment complex structures even when they are incomplete or fragmented. By integrating PartSAM with the part completion model HoloPart (Yang et al., 2025), these segmented fragments can be seamlessly reconstructed, achieving a comprehensive and complete decomposition of the mesh. This combined approach ensures that the segmented parts are not only accurate but also fully realized, overcoming the limitations of standard segmentation techniques.
@@ A.2.4 Application
A.2.5 Results on PartObjaverse-Tiny with Mesh Connectivity
In the evaluation of automatic segmentation on PartObjaverse-Tiny (Section 4.2), we compare PartField (Liu et al., 2025) and our method in a setting where mesh connectivity is disabled. The rationale behind this choice is further explained through the quantitative results in Table 4. Rather than training PartField, we use a randomly initialized model to generate the feature field and apply agglomerative clustering to produce segmentation results with mesh connectivity (PartField (B) in Table 4). Interestingly, the segmentation performance is significantly better than that of the pre-trained PartField when mesh connectivity is disabled (PartField (A)), with a $27\%$ improvement.
@@ A.2.5 Results on PartObjaverse-Tiny with Mesh Connectivity
Moreover, as indicated by the results in the last row of Table 4, we find that PartSAM can also benefit from mesh connectivity information. After getting the per-face segmentation labels using our segment-every-part mode, we apply graph cuts to refine the segmentation, following Boykov et al. (2001). Specifically, we construct a mesh graph where each face is treated as a node, and edges represent adjacency relationships between faces. The graph cut algorithm minimizes an energy function that balances two key components: (1) a data term which measures the disagreement between the refined label assignment and the initial per-face predictions, and (2) a smoothness term that penalizes large label changes between adjacent faces, encouraging spatially coherent boundaries. By minimizing this energy function, this post-processing ensures adjacent faces have consistent labels, thereby improving segmentation accuracy. By using connectivity information in this way, PartSAM outperforms PartField under this setting (Ours (B)), further demonstrating the flexibility of our method.
A.2.6 Complexity Analysis
@@
We present the complexity analysis in Table 5, reporting both the inference time on a single NVIDIA H20 GPU and the number of network parameters. Compared with lifting-based methods that require per-shape optimization (Tang et al., 2024; Yang et al., 2024b), PartSAM achieves substantially faster inference due to its feed-forward design. Compared with PartField (Liu et al., 2025), PartSAM incurs only a slight increase in inference time for segmenting an entire shape. Although our encoder is built upon PartField, it eliminates the need for per-face dense feature sampling and clustering. Furthermore, our decoder is lightweight, ensuring that the overall inference time remains well-controlled. In addition, we further report the runtime and performance under different numbers of prompt points (1024/512/256). The results show that PartSAM delivers consistently strong performance across all settings, indicating that our method is robust to the choice of prompt density. These results demonstrate that our method effectively balances computational cost and performance.
A.2.7 Cross-shape Consistency.
We follow the evaluation protocol of Liu et al. (2025) and use functional maps (Ovsjanikov et al., 2012) to compute correspondences between pairs of shapes in Figure 13. For similar quadruped shapes (first row), both PartField and our PartSAM achieve reasonable cross-shape consistency, yet ours yields clearer correspondences in fine-grained regions such as the tail. When the structural gap becomes larger (second row), our method remains robust—producing correct correspondences on challenging parts like the hands—whereas PartField fails and yields entirely incorrect mappings. These results highlight the stronger part-aware consistency of our learned features under large-scale native 3D supervision.
A.2.8 Failure Cases and Limitations.
We present failure cases of PartSAM in Figure 14, and provide additional results on diverse meshes in Figures 17 and 18, where failure cases are highlighted using red dashed boxes. Although PartSAM benefits from large-scale native 3D supervision, the diversity of existing 3D datasets (Deitke et al., 2023b; a) remains limited compared with high-coverage 2D datasets such as SA-1B (Kirillov et al., 2023). Consequently, structures that rarely appear during curation may not be accurately segmented at test time. For example, in the first row of Figure 14, PartSAM fails to extract the engraved letters carved into surfaces—these surface-level markings never appear as valid parts in current 3D pipelines. SAMesh can sometimes detect such tiny engraved regions due to its direct use of SAM outputs, though its multi-view masks may be inconsistent and introduce noise. PartField, which relies on feature space clustering, is even less able to recover such fine-grained details. Another type of failure arises when an object lacks a clear semantic structure, as exemplified by the coral sculpture in the second row. In these cases, both PartSAM and existing baselines struggle to produce meaningful decompositions because the object itself does not possess a coherent part hierarchy.
Figure 13: Point-to-point correspondences obtained by Functional Maps (Ovsjanikov et al., 2012) using features of PartField (Liu et al., 2025) and our encoder as input.
These cases highlight a limitation of current 3D data curation: the class-agnostic nature of part annotation, combined with the inherent ambiguity of open-world part definitions, leads to a long-tail distribution where many part types appear only rarely. Such an imbalance makes it difficult for the model to distinguish the rare, minority parts. This issue becomes even more pronounced on AI-generated meshes or coarse real-world scans, where geometric irregularities, incomplete structures, or hallucinated details further amplify the ambiguity of part boundaries and exacerbate the difficulty of consistent part decomposition.
Meanwhile, although PartSAM achieves state-of-the-art performance in class-agnostic part segmentation, similar to 2D SAM (Kirillov et al., 2023; Ravi et al., ), it cannot directly produce semantic labels for these masks, which are important for some downstream tasks. One possible future direction to address this limitation is to generate large-scale training datasets consisting of 3D shapes, part annotations, and corresponding semantic labels. This could potentially be achieved by utilizing PartSAM’s interactive segmentation capabilities, where users provide feedback on segmented parts, facilitating the assignment of semantic categories.
@@
Methods Time Params Performance
SAMesh $\sim$ 7min / 56.9
SAMPart3D $\sim$ 15min 114M 53.5
PartField $\sim$ 10s 106M 51.5
[…]
$\sim$ 9s (768 points) 68.8
$\sim$ 7s (512 points) 67.9
$\sim$ 4s (256 points) 65.2
A.3 Evaluation on 3DCoMPaT++
To comprehensively evaluate PartSAM, we compare it with PartField (Liu et al., 2025) and our ablation variant trained without the second-stage native 3D data on a recent dataset, 3DCoMPaT++ (Slim et al., 2025). Since the part annotations in 3DCoMPaT++ are semantic-level labels, both the baselines and PartSAM experience performance drops compared with the results on instance-level datasets. Nevertheless, PartSAM still significantly outperforms PartField and also surpasses our ablation variant trained without the second-stage native 3D supervision, highlighting the importance of large-scale 3D data in achieving strong generalization. This performance gap aligns with the characteristics of the dataset. 3DCoMPaT++ contains extremely fine-grained part annotations, which pose challenges for clustering-based methods like PartField—whose clustering step struggles to reliably separate small or rare parts. In contrast, our promptable decoder naturally produces masks across multiple granularities, enabling the model to recognize and segment diverse structures within a unified framework. This allows PartSAM to handle the fine-grained decomposition required in 3DCoMPaT++.
[…]
Figure 14: Failure cases of PartSAM.
A.4 Additional Visualization
@@
Visualization of Segmentation Results Additional segmentation results of PartSAM are shown in Figure 16, where it consistently delivers impressive performance across different kinds of shapes. To provide a more comprehensive picture of PartSAM’s behavior, we also present randomly sampled results in Figure 17 and Figure 18, where failure cases are highlighted with red dashed boxes.
Table 6: Quantitative comparison of automatic segmentation on 3DCoMPAT++ (Slim et al., 2025). The best scores are emphasized in bold. We report the mean IoU.
Category Method Coarse Fine-grained Avg
Plane PartField 47.3 30.5 38.9
Ours w/o Model-in-the-loop Annotation 57.7 42.6 50.2
Ours 59.4 47.5 53.5
Car PartField 40.5 20.7 30.6
Ours w/o Model-in-the-loop Annotation 54.1 33.5 43.8
Ours 57.9 37.7 47.8
Bag PartField 51.9 41.4 46.7
Ours w/o Model-in-the-loop Annotation 61.2 52.7 57.0
Ours 66.9 55.3 61.1
Coat Rack PartField 48.0 44.7 46.4
Ours w/o Model-in-the-loop Annotation 58.8 46.3 52.6
Ours 62.5 50.2 56.4
Toilet PartField 44.2 38.7 41.5
Ours w/o Model-in-the-loop Annotation 58.4 62.2 60.3
Ours 60.3 65.7 63.0
@@
[…]
Figure 17: Randomly sampled automatic segmentation results on AI-generated shapes (Lai et al., 2025). Failure cases are highlighted with red dashed boxes.
[…]
Figure 18: Randomly sampled automatic segmentation results on Yang et al. (2024b) and real-world scans (Choi et al., 2016; Wu et al., 2023). Failure cases are highlighted with red dashed boxes.